With some services, e.g., checking accounts, phone service, or pay TV, a consumer is offered a choice of two or more payment plans. One can either pay a high entry fee and get a low price per unit of service or pay a low entry fee and a high price per unit of service. Suppose you have an income of $100. There are two plans. Plan A has an entry fee of $20 with a price of $2 per unit. Plan B has an entry fee of $40 with a price of $1 per unit for using the service. Let x be expenditure on other goods and y be consumption of the service.

a. Write down the budget equation that you would have after you paid the entry fee for each of the two plans.

b. If your utility function is xy, how much y would you choose in each case?

c. Which plan would you prefer? Explain.
Answer #1

- Answers Page : 01 Given that plan A his an entry fee of 420 with a poice of $2 per unit. And plan B has an entry fee of & tPage : 02 SO, MRS = MUX / MUY In optamom condition , MRS= Pxley (8) (11x): 2/1 = 2 y : 2x X Substeluiting in budget keine forx=20 Y = 2x - 2420= 40. ④ Total stakty: Given of otality fonction is XY. for pho A : XY = 16832 = 512 for plan B : XY 20x40 =